OpenStudy (anonymous):

On the first day Louis read half of the book, on the second day he read 3/8 more and 40 pages remain. How may pages where their in the book?

OpenStudy (michele_laino):

Pleae, note that, your problem is a mathematization problem. I think that If I call n, the number of pages of your book, we can write: \[\frac{ n }{ 2 }+\frac{ 3 }{ 8 }n+40=n\] because on the first day Louis read n/2 pages, on the secon day, Louis read (3/8)*n more and 40 pages reamain, so the sum below: n/2 + (3/8)*n+40, must be equal to n. Please try to solve the above equation in order to find the value of n

OpenStudy (anonymous):

what is the answer?

OpenStudy (michele_laino):

Sorry, but by the code of conduct, I can not tell you the answer directly!

OpenStudy (anonymous):

how to get the answer? please help me. i need it right now

OpenStudy (michele_laino):

you have to solve the above equation for n, in order to that, you can calculate the least common multiple. Do you know what is the least common multiple among 1, 8 and 2?

OpenStudy (anonymous):

no

OpenStudy (michele_laino):

do you know the rule to apply to calculate the least common multiple?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I think its 16

OpenStudy (michele_laino):

ok! now I write the subsequent factorizations: 1=1 2=1*2, 8=2*2*2=2^3 please note that in order to find the least common multiple (LCM) you have to pick common factors with highest exponent, and all not common factors with highest exponent, so?

OpenStudy (anonymous):

16

OpenStudy (michele_laino):

Sorry, what is the highest exponent of 2?

OpenStudy (anonymous):

3

OpenStudy (michele_laino):

so, LCM=...

OpenStudy (anonymous):

8

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

what will i do next?

OpenStudy (michele_laino):

I write. Next step you have to multiply both sides of the equation that I wrote, by 8, please write what you get

OpenStudy (anonymous):

what is the answer i will try to figure out?

OpenStudy (michele_laino):

for example the left side becoes this: \[8*\frac{ n }{ 2 }+8*\frac{ 3 }{ 8 }n+8*40=\] \[=4*n+3*n+320=\] \[=7*n+320\] now, do the same with the right side, please!

OpenStudy (anonymous):

n=320

OpenStudy (michele_laino):

that's right, nevertheless I ask you, what is the right side of the equation?

OpenStudy (anonymous):

8n

OpenStudy (anonymous):

so waht is the answer?

OpenStudy (michele_laino):

that's right, so we have to solve the subsequent equation: \[7*n+320=8*n\] now, in order to that, please add (-7*n) to both sides of that last equation

OpenStudy (anonymous):

n=320

OpenStudy (anonymous):

then? what will i do next how will i know the pages of the book?

OpenStudy (michele_laino):

ok!we have: n=20, so the book of Louis has 320 pages!

OpenStudy (michele_laino):

sorry n=320....

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how to check the answer?

OpenStudy (michele_laino):

by the correctness of our reasoning

OpenStudy (anonymous):

how?

OpenStudy (michele_laino):

I think that our reasoning is correct, so our answer is right, that's all! you have to trust of your reasoning!

OpenStudy (anonymous):

ok btw, Thanks again,

OpenStudy (michele_laino):

thank you!